Status Report

NASA Space Station On-Orbit Status 19 Feb 2004

By SpaceRef Editor
February 19, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 19 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.

The crew successfully completed the Orlan/Soyuz ingress demo as planned, and everything went very well.   [Actual transition from the DC-1 “Pirs” to the Soyuz orbital module (BO) in pressurized and tethered suits lasted just about 21 minutes.  Suit donning and setting up coolant flow took about 20 minutes more than estimated, but the crew generally was able to stay on the timeline.   The demo started last night at ~11:15pm EST with teardown & removal of the air ducts between the DC-1 and BO, followed by Orlan ingress, checkout of comm & biomedical telemetry via the BSS interface system for vital signs and equipment monitoring, final tests of Soyuz TMA-3 systems, functionality checkout of suit and hatch valves, etc.  Foale had to re-adjust sleeve length on his Orlan-M #14.  Actual translation began at ~5:00am, and both crewmembers were able to ingress the Soyuz BO quickly without real issues, followed by hatch closing and Orlan (“Eagle”) doffing inside the BO.]

After exercise completion, the crew reinstalled the air ducts and reset Soyuz and DC-1 thermal control systems (SOTR) to their original local temperature commutators (LKT) configuration, which deactivated the “Transit” EVA radio telemetry (T/M) system and switched from DC-1 T/M back to Soyuz T/M. 

FE Alex Kaleri downloaded data and imagery collected yesterday of the BIO-5 Rasteniya-2 (“Plants-2”) experiment to the computer for subsequent downlink to the ground.   [Rasteniya studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse.]

CDR/SO Mike Foale worked on the smoothly-performing MSG (Microgravity Science Glovebox) for his regular support of the on-going PromISS-3 (Protein Crystal Growth Monitoring by Digital Holographic Microscope 3) experiment, today removing videotape #9 and installing tape #10.

Mike also conducted the daily routine maintenance on the SOZh life support system, comprising the water supply equipment, food supply subsystem (SOP), and sanitary hygiene equipment (SGO).

The crew worked out according to their regular daily physical exercise program on TVIS treadmill and CEVIS cycle (aerobic), VELO ergometer with force loader and RED (anaerobic).

At dinnertime (10:30am), as every day, the crew supported the Renal Stone prevention experiment by taking the test medication (either potassium citrate or placebo tablets) until the next sample collection phase in early April this year.

Yesterday’s ground-commanded transition of the U.S. segment’s (USOS) internal thermal control system (ITCS) from single LTL (low temperature loop) to dual-loop mode was successfully performed, in preparation to support the EVA-9 next week.   [The transition included a resetting of the temperature setpoint of the MTL TWMV (moderate temperature loop/three-way mixing valve) from 12.2 degC to 17.2 degC.  The TWMV controllers were observed to respond somewhat slower than expected to regulate outflow temperature back down when it exceeded the setpoint.  The issue, which has had no negative hardware impacts, is under study.]

After the crew had doffed their EVA gear, Russian TCS (thermal control system) was switched from the KOV-2 loop to KOV-1 on replaceable pump panel #1 (SPN-1).  [This configuration gives rise to less than optimal condensate collection for the SKV-1 dehumidifier and increases the condensate amount collected in the quickly filling-up Lab condensate collection tank.]

Also preparatory to the EVA-9, MCC-H conducted a test in real time of the ability to send commands and time tags to the USOS through Russian assets.  The command transfer was successful, but the time tag transmission did not go beyond MCC-Moscow.  Teams are reviewing the cause.    [Purpose of test was (a) to provide proficiency to ground controllers in the event that this contingency comm configuration is required during the spacewalk, and (b) to acquire data on command-link “latency”, i.e., the time lag between command uplink and command execution, given that in a worst-case situation the command has to “queue up” on the RS server, which may be loaded with interfering Russian signals.]

Comparison O2 readings taken yesterday with on-board CSA-CP (compound specific analyzer-combustion products) instruments, the MCA (major constituent analyzer) and the SM GA (gas analyzer) again showed an inaccurate reading by CSA-CP #1009 (O2 = 16.3%), compared to the MCA (21.86%), the GA (19.81%) and the primary CSA-CP #1003 (21.8%), one of the two new units delivered on 13P.   [Moscow has requested a calibration of the GA against the MCA tomorrow (2/20), because of its accuracy.]

Relative motion (RELMO) analysis by MCC-H on the Progress solar-array restraint bolt that drifted away from the ISS on 2/15 shows that the atmospheric drag on the bolt is much higher than originally assumed on the day of the event.  Therefore, the bolt has left the vicinity of the station very quickly, and any re-contact risk to the ISS, extremely low to start with, has disappeared.   [MCC-M described the bolt as having 36 grams mass, a drag coefficient of 2.35 and an effective drag area of ~830 square millimeters.]

The crew’s sleep time began at 12:30pm EST.  Tonight they can sleep two hours longer (to 11:00pm) than yesterday, with bedtime also slipping to two hours later (2:30pm).

Today’s CEO targets, excluding almost all targets in the Americas due to dominantly descending orbit tracks combined with the sleep shift, were Jarvis Island, Pacific (400mm-lens.  This two-mile long equatorial atoll is completely surrounded by a coral reef), Beijing, China (ideal near-vertical pass.  Looking slightly left of nadir for the city center), Tianjin, China (Beijing’s seaport, slightly left of nadir.  Beijing and Tianjin are growing towards one another [like Houston and Galveston].  The extent of land use change in this corridor is of great interest, as well as images of the extent of these cities), Yellow River Delta, China (the mouth of the Yellow has changed dramatically even during the period of manned spaceflight, partly due to the extreme silt load it carries, and partly due to human intervention), Internal waves, Vietnam (looking right near the early afternoon glint point for probable internal waves generated off central Vietnam), Guangzhou, China (pass along the west side of this major urban region on the banks of the Pearl River), Internal waves, Philippines (looking right towards the glint disc for water flow patterns and internal waves set up around the many islands that stretch between Mindanao [Philippine’s large southern island] and the northern tip of Borneo), Calcutta, India (nadir pass over this city of 13 million, stretched out on both sides of the westernmost tributary of the Ganges delta), and Betsiboka delta, Madagascar (400mm-lens.  Nadir pass.  Researchers have an excellent time series of the evolution of sedimentation in this estuary on NW Madagascar.  The crew’s recent imagery shows a sudden increase in the pace of change: most seaward islands growing by the addition of sediment, the emergence of two new islands, and the narrowing of waterways in the network.  Detailed images are requested to help understand how much of the change is related to tidal change and how much to absolute island growth.  Deforestation inland since World War II is blamed for the massive downstream movement of soil as river sediment).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 3:00am EST, 2/13).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered Off.  Vozdukh CO2 scrubber is On (in Automatic Mode).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (repair now completed; to be tested ASAP).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.1; ppO2 (mmHg) — 145.9; ppCO2 (mmHg) — 2.5;
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.9.
  • FGB Cabin:  Pressure (mmHg) — 744; temperature (deg C) — 22.7.
  • Node:  Pressure (mmHg) — 740.71; temperature (deg C) — 23.6 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 742.59; temperature (deg C) — 24.8; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 742.80; temperature (deg C) — 23.7; shell heater temp (deg C) — 24.1, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 22.5
  • PMA-2:  Shell heater temp (deg C) — 12.1

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed Position (2B: 235 deg; 4B: 125 deg); non-suntracking, “night glider”/”sun slicer” drag reduction mode.
  • SM batteries:  Battery #8, formerly known as #7, is failed (to be replaced); all other batteries (7) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #3 is off (capacity restoration mode, ROM); battery #6 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 4070 kg (8972 lb) as of 2/12/04  [SM(755) + FGB(2656) + Progress M(0) + Progress M-1(659)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • *S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 6:49am EST [= epoch]):

  • Mean altitude — 366.7 km
  • Apogee — 371.2km
  • Perigee — 362.2 km
  • Period — 91.88 min.
  • Inclination (to Equator) —  51.628 deg
  • Eccentricity — 0.0006682
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 90 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29969

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.